### v3

# PVChat: Personalized Video Chat with One-Shot Learning

Abstract
Video large language models (ViLLMs) excel in general video understanding, e.g., recognizing activities like talking and eating, but struggle with identity-aware comprehension, such as “Wilson is receiving chemotherapy” or “Tom is discussing with Sarah”, limiting their applicability in smart healthcare and smart home environments. To address this limitation, we propose a one-shot learning framework PVChat, the first personalized ViLLM that enables subject-aware question answering (QA) from a single video for each subject. Our approach optimizes a Mixture-of-Heads (MoH) enhanced ViLLM on a synthetically augmented video-QA dataset, leveraging a progressive image-to-video learning strategy. Specifically, we introduce an automated augmentation pipeline that synthesizes identity-preserving positive samples and retrieves hard negatives from existing video corpora, generating a diverse training dataset with four QA types: existence, appearance, action, and location inquiries. To enhance subject-specific learning, we propose a ReLU Routing MoH attention mechanism, alongside two novel objectives: (1) Smooth Proximity Regularization for progressive learning through exponential distance scaling and (2) Head Activation Enhancement for balanced attention routing. Finally, we adopt a two-stage training strategy, transitioning from image pre-training to video fine-tuning, enabling a gradual learning process from static attributes to dynamic representations. We evaluate PVChat on diverse datasets covering medical scenarios, TV series, anime, and real-world footage, demonstrating its superiority in personalized feature understanding after learning from a single video, compared to state-of-the-art ViLLMs.
Figure 1: Examples of PVChat’s ability with one-shot learning (e.g., <Nz>and <Ab>). PVChat can answer questions about the personalized information correctly while other models [5, 50] fail.
1 Introduction
Recent advancements in Video Large Language Models (ViLLMs) have showcased impressive capabilities across a wide range of tasks, e.g., video captioning, temporal action localization, and video question answering [5, 50, 31, 49]. These models, trained on vast datasets, exhibit broad knowledge across multiple domains, such as medical assistance [17] and autonomous driving [55], significantly extending the functional scope of ViLLMs.
Despite these achievements, current ViLLMs are largely confined to general-purpose content understanding and struggle in personalized scenarios that require distinguishing specific individuals. For instance, when analyzing a video featuring a person named “Alex”, even state-of-the-art models fail to consistently recognize “Alex” in novel contexts, such as identifying him from others or describing his specific actions. This limitation arises because existing training data predominantly focus on general knowledge rather than learning identity-specific information about an individual. As a result, ViLLMs are ill-equipped for real-world applications that demand personalized comprehension, e.g., smart home [60], intelligent healthcare [15] and human-computer interaction [6].
To enhance the ability of LLMs in capturing personalized cues, recent research has focused on personalized image understanding [34, 38, 37]. For example, Yo-LLaVA [34] acquires the capability to encode a customized subject into a collection of latent tokens by utilizing a small number of sample images representing the subject. Other studies [38, 37] construct specialized training datasets curated with expert annotations to enable personalized image recognition. While these methods have demonstrated promising results, they are inherently limited to static image-based understanding and cannot model dynamic, temporally evolving visual cues present in videos. Given that videos contain significantly richer personalized information, including motion patterns, interaction dynamics, and contextual dependencies, there is a pressing need for solutions that enable identity-aware comprehension in video-based contexts.
To address these limitations, we propose PVChat, the first personalized video understanding LLM capable of learning individual characteristics and supporting subject-specific question answering from a single reference video. To enable one-shot learning, we propose a data augmentation pipeline that synthesizes high-quality personalized samples. It generates identity-preserving positives while incorporating visually similar negatives for enhanced discrimination. The process begins with facial attribute extraction and demographic categorization, followed by high-fidelity video synthesis using ConsisID [56] and PhotoMaker [21]. Hard negatives are introduced by retrieving similar faces from Laion-Face-5B [59] and CelebV-HQ [61] with corresponding video synthesis. Finally, question-answer pairs covering four key categories—subject existence, appearance, action, and location—are generated via InternVideo2 [50] and ChatGPT-4o to ensure linguistic coherence and personalization.
To improve subject-specific feature learning, we introduce a ReLU Routing Mixture-of-Heads (ReMoH) attention mechanism, which replaces conventional softmax-based and top-k head selection with a ReLU-driven dynamic routing strategy, enabling smooth and scalable training. Additionally, we propose two novel optimization objectives: Smooth Proximity Regularization, which promotes progressive learning via exponential distance scaling, and Head Activation Enhancement, which balances shared and routed attention heads, effectively mitigating gradient explosion and inactive heads in multi-head attention mechanisms. To optimize our ReMoH-inspired PVChat under limited data conditions, we adopt a two-stage training strategy transitioning from static image learning to video temporal modeling. In the first stage, the model is trained on images with a simple summary task to capture static identity attributes. In the second stage, it is fine-tuned on video QA tasks to develop dynamic action recognition and multi-person interaction capabilities. This progressive learning framework enables a structured transition from static representations to complex spatiotemporal reasoning. We verify PVChat across a diverse range of personalized scenarios, including healthcare, TV series, anime, and real-world scenes, requiring the recognition of one, two, or three individuals. Experimental results demonstrate that PVChat achieves state-of-the-art performance in individual information comprehension and identity-aware reasoning from a single reference video.
In summary, our main contributions of this work are listed as follows:
  • We introduce PVChat, the first ViLLM that extends LLMs with personalization capability from a single reference video. This enables personalized video understanding and user-specific question-answering with only one-shot learning, laying a strong foundation for future individual video comprehension.
  • We develop a systematic video augmentation and QA generation pipeline from a single reference video. The key is to exploit off-the-shelf toolboxes to generate identity-preserving positives and synthesize hard negatives for confusingly similar faces. To support the research, we construct a diverse dataset containing $6$ individual scenarios, $304$ original videos, $2{,}304$ extended generated videos with over $30{,}000$ QA pairs, which will be publicly released to facilitate future research.
  • We design the ReLU Routing Mixture-of-Heads module to enhance the effective extraction of individual-specific characteristics from videos. Additionally, we introduce Smooth Proximity Regularization and Head Activation Enhancement to improve training stability and attention head activation.
2 Related Works
2.1 Image Large Language Models
The remarkable language comprehension and reasoning capabilities of Large Language Models (LLMs) [1, 3, 45, 47] have spurred significant interest in extending them into multimodal understanding. Pioneering works demonstrate different approaches for vision-language alignment: Flamingo [47] processes interleaved image-text inputs through cross-attention layers for diverse multimodal tasks, while BLIP-2 [16] employs a Querying Transformer (Q-Former) to bridge frozen visual encoders with LLMs. Simpler architecture like LLaVA [24] achieves effective feature alignment using lightweight MLP projectors. Recent research systematically explores critical components of Multimodal LLMs (MLLMs), including dynamic high-resolution processing strategies [4, 26], instruction-tuning methodologies [18, 28], and optimal visual encoder selection [46, 52]. Comprehensive design space analyses from MM1 [32] and Idefics2 [14] further establish foundational principles for MLLM development.
2.2 Video Large Language Models
With progress in image-based multimodal language models (MLLMs), video understanding research has garnered great attention. Video-ChatGPT [30] and Vally [29] aggregate temporal features into compact tokens, while Video-LLaVA [22] unifies image-video representations through aligned projections before LLM integration. Video-Teller [23] underscores modality alignment in pre-training objectives, whereas PLLaVA [54] examines feature pooling impacts on video Question-Answering tasks. For extended sequences, LLaMA-VID [19] encodes frames with dual tokens to balance efficiency and information retention, while MovieChat [44] employs memory optimization for processing ultra-long videos (>10K frames). Weng et al. [53] enhance local segment features by injecting global semantics through hierarchical fusion mechanisms. These approaches collectively address spatiotemporal modeling challenges across varied input scales. However, these models are still incapable of understanding videos with personalization.
2.3 Personalized Large Language Models
Personalization in generative models exhibits distinct implementations across domains. For image generation, existing approaches focus on pixel-level subject fidelity through either concept token optimization [8, 13] or model parameter adaptation [40]. In natural language processing contexts, personalization typically involves shaping LLMs’ behavioral patterns (e.g., , conversational style) via prompt engineering or metadata-driven retrieval mechanisms [27, 58, 35]. Recent personalized modeling approaches employ parameter-efficient adaptation through few-shot learning. MyVLM [2] implements a Q-former-based architecture with trainable projection heads for concept extraction and visual feature augmentation, whereas Yo’LLaVA [33] extends the LLaVA framework by integrating novel concepts as specialized tokens within the LLM’s embedding space. To the best of our knowledge, our model is the first to support videos input while others only accept images as inputs.
3 PVChat
3.1 Overview
To achieve video personalization understanding, we first establish a promising and systematic data augmentation pipeline in 3.2 to deal with the severe lack of diverse video data for individuals. To enhance the subject-specific learning, we propose a ReLU Routing MoH Attention mechanism in 3.3, alongside two novel objectives: (1) Smooth Proximity Regularization (SPR) for progressive learning through exponential distance scaling; (2) Head Activation Enhancement (HAE) for balanced attention routing. Finally, a two-stage training strategy is adopted in 3.4 to progressively learn from static features to dynamic features of the target.
3.2 Data Collection
Figure 2: The systematic data collection pipeline. For positive data collection, the original videos are processed by DeepFaceLab [36] for high-quality face and InterVideo2 [50] for demographic characteristics, which boost identity preservation. ConsisID [56] and LivePortrait [9] with PhotoMaker [21] utilize the identity information to generate videos of various background or different motion/expression, respectively. For model’s robust perception, hard negative samples are selected from either similar face retrieval to generate negative videos, or sampled from the CelebV-HQ dataset [61]. These negative samples guarantee the model’s accurate recognition of both identity and content.
To address the challenge of limited training data and achieve one-shot learning, we propose a novel and comprehensive data augmentation pipeline as illustrated in Fig.2, which starts from two key principles: 1) enabling automated personalized data collection directly from a wide range of natural videos; 2) ensuring that the generated personalized data encompasses different scenes, actions, and contextual elements while preserving the identity information.
Identity-Preserving Positives Generation. Maintaining the identity information is key to obtaining diverse data of a specific individual. To achieve that, our pipeline starts with facial extraction from original videos using DeepFaceLab [36]. When multiple faces are present, we carefully design a two-stage approach: first, we extract facial embeddings via FaceNet [41], then apply DBSCAN [7] clustering to differentiate between subjects. After that, the faces are evaluated using several metrics, including Eye Aspect Ratio (EAR) to determine eye openness (with a threshold of 0.25), facial orientation, image clarity, and presence of facial landmarks. These metrics allow us to identify the highest-quality facial image (HQ-face) for subsequent steps.
To further enhance the stability of identity preservation, the characteristics of individuals in the videos are determined using InternVideo2 [50], which are categorized by gender (male/female) and age group (young/middle-aged/elderly) based on the video content and our designed prompt. With these demographic characteristics, we can more accurately preserve the identity, and employ ConsisID [56] with our extensive prompt library Fig. 6 to generate a wide range of videos in different scenarios (e.g., gym, restaurant, school, etc) with identity consistency. Although videos from ConsisID [56] contain rich content and we extract multiple types of identity information, their ID consistency is still poor, leading to a decline in data quality.
With the aim of incorporating more ID consistent data, we utilize PhotoMaker [21] with the identified characteristics to generate additional high-fidelity images of the subject across different environments, and adopt LivePortrait [9] to animate these images to generate high-identity-consistency videos with a facial motion and expression library (including nodding, head-shaking, questioning, anger, and smiling). Meanwhile, we also directly input the HQ-face image to LivePortrait [9], generating the motion video. These videos exhibit strong ID consistency but include simple contents, complementing the data from ConsisID [56].
Hard Negatives Retrieval. Avoiding the misidentification of personalized subjects is crucial for personalized video chatting. As observed in previous VLM research [34] and our video chat experiments, training with only positive data can lead to the model losing its ability to recognize personalized subjects. Considering these factors, we introduce challenging negative samples by retrieving top-k visually similar faces from the Laion-face-5b dataset [59] utilizing CLIP [39]. Similarly, these facial images are animated by LivePortrait [9] to serve as negative data, which enables the model to learn to robustly identify the personalized subjects. In addition, these negative samples also contain relatively simple movements and other content, which inspires us to supplement them with 30 randomly selected videos from the CelebV-HQ dataset [61] that possesses rich and high-resolution content. Combining them together, the personalized video chat model is equipped with robustness to both subjects and content.
Question-Answer Pairs Generation. Having various positive and negative samples, the final stage involves generating question-answer pairs using InternVideo2 [50] across four categories: existence verification, appearance description, location identification, and action recognition. To enhance the linguistic naturalness, we refine these responses using ChatGPT-4o, converting generic human references into subject-specific references. The QA generation is illustrated in Fig. 3.
Through this comprehensive pipeline, each input natural video is transformed into $81$ different videos accompanied by $1{,}455$ question-answer pairs, significantly expanding our training dataset while maintaining subject identity consistency and guaranteeing the model’s robustness.
3.3 ReLU Routing Mixture-of-Heads Attention
Multi-Head attention. We begin by reviewing the standard multi-head attention mechanism [48] which is based on scaled dot-product attention. For $T_{1}$ input tokens $\boldsymbol{X}_{1}\in\mathbb{R}^{T_{1}\times d_{in}}$ where $d_{in}$ is the input dimension, and $T_{2}$ tokens $\boldsymbol{X}_{2}\in\mathbb{R}^{T_{2}\times d_{in}}$, attention is conducted as follows:
[…]
To enhance the representation of the attention layer, Vaswani [48] proposed to use multiple attention heads to operate on the projection of the input. The transformer computes $h$ different projection of $(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})$, and the concatenation form of multi-head attention is
$\begin{array}[]{c}\operatorname{MHA}\left(\boldsymbol{X}_{1},\boldsymbol{X}_{2}\right)=\text{ Concat }\left(\boldsymbol{H}^{1},\boldsymbol{H}^{2},\ldots,\boldsymbol{H}^{h}\right)\boldsymbol{W}_{O},\\ \boldsymbol{H}^{i}=\operatorname{Attention}\left(\boldsymbol{X}_{1}\boldsymbol{W}_{Q}^{i},\boldsymbol{X}_{2}\boldsymbol{W}_{K}^{i},\boldsymbol{X}_{2}\boldsymbol{W}_{V}^{i}\right),\end{array}$ (2)
where $\boldsymbol{W}_{Q}^{i},\boldsymbol{W}_{K}^{i},\boldsymbol{W}_{V}^{i}$ represent the $i_{th}$ projection matrices of the query, key and value. $\boldsymbol{W}_{O}$ is the final projection matrix, which is decomposed by $\boldsymbol{W}_{O}=\left[\boldsymbol{W}_{O}^{1},\boldsymbol{W}_{O}^{2},\ldots,\boldsymbol{W}_{O}^{h}\right]$. The summation form is as follows:
$\operatorname{MHA}\left(\boldsymbol{X}_{1},\boldsymbol{X}_{2}\right)=\sum_{i=1}^{h}\boldsymbol{H}^{i}\boldsymbol{W}_{O}^{i}.$ (3)
Figure 3: We illustrate the process of automatically generating question-answer pairs using InternVideo2 [50] and ChatGPT [1]. A positive and a negative sample are shown at the bottom.
3.3.1 Heads as Experts with ReLU Routing
Motivation of ReMoH. Recent study Mixture-of-Heads [12] targets to allow each token to select the Top-k relevant heads, improving inference efficiency without sacrificing accuracy or increasing the parameters, compared with traditional multi-head attention. Moreover, MoH [12] finds it can learn domain-specific information, which is consistent with our personalized video understanding with only one-shot learning. However, the vanilla Top-k choice of MoH limits the performance and domain-specific information learning, which is because: 1) Top-k choice is not fully differentiable for training; 2) The choice of heads is not adaptive and flexible, and interferes with the domain-specific knowledge learned by expert heads. Thus, inspired by [51], our ReLU Routing Mixture-of-Heads Attention (ReMoH) aims at enhancing MoH with much more efficient and adaptive learning especially for the personalized subject’s information, with only an increase of 2 MLP parameters. The detailed design for ReMoH can be seen in Fig. 4 (b).
ReMoH with Head and ReLU Routers. Specifically, ReMoH consists of $h$ attention heads $\boldsymbol{H}=\left\{H^{1},H^{2},\ldots,H^{h}\right\}$ and uses a head routers with a ReLU router to modulate the output of different expert heads. Formally, given input tokens $\boldsymbol{X}_{1}$ and $\boldsymbol{X}_{2}$, the output of ReMoH is the weighted sum of outputs of the heads:
$\begin{array}[]{c}\operatorname{ReMoH}\left(\boldsymbol{X}_{1},\boldsymbol{X}_{2}\right)=\sum_{i=1}^{h}s_{i}\boldsymbol{H}^{i}\boldsymbol{W}_{O}^{i},\end{array}$ (4)
where $s_{i}$ is the score corresponding to $H^{i}$. In the attention mechanism, some attention heads capture common knowledge across different contexts, such as grammatical rules in language, while others focus on context-specific knowledge [42]. Following MoH [12], we divide the heads into shared heads and routed heads, where shared heads are always activated, and routed heads will only be activated if the ReLU routing score is non-zero. Specifically, the score for each head is defined as:
$s_{i}=\left\{\begin{array}[]{ll}\alpha_{1},&\text{ if }1\leq i\leq n,\\ \alpha_{2}\operatorname{ReLU}\left(\boldsymbol{W}_{r}\boldsymbol{x}_{t}\right)_{i},&\text{ if }n<i\leq n+m,\end{array}\right.$ (5)
where $n$ and $m$ are the number of shared heads and routed heads, respectively. $\boldsymbol{W}_{r}\in\mathbb{R}^{m\times d_{in}}$ denotes the projection matrix of the ReLU router. $\alpha_{1}$ and $\alpha_{2}$, which come from the shared router, will be used to balance the contributions of the shared and routed heads, and are defined as:
$\left[\alpha_{1},\alpha_{2}\right]=\operatorname{Softmax}\left(\boldsymbol{W}_{h}\boldsymbol{x}_{t}\right).$ (6)
With the utility of ReLU router to choose specific routed heads to be activated, the decision process becomes fully differentiable and offers greater flexibility. This boosts the learning of the personalized video information, avoiding a fixed number of activated experts to limit the performance or cause computational redundancy. A more detailed analysis can be found in our ablation study.
Figure 4: (a) The training pipeline of our method. (b) The proposed ReMoH technique for better specialized characteristics learning.
3.3.2 Controlling Sparsity Strategy
One important issue found in [51] is that most experts are always activated during training. Our ReMoH aims to control computation cost by managing the sparsity of the ReLU output, targeting a sparsity of $T_{s}=(1-\frac{b}{m})$, where $b$ is the number of routed heads we set for encouraging the activation. To encourage the sparsity of activation, which helps to reduce computational cost and boost specific-domain learning, we propose the Smooth Proximity Regularization (SPR) Loss to make the training more flexible:
$\mathcal{L}_{SPR}=\beta_{p}\cdot\mathcal{L}_{Reg},$ (7)
where $\beta_{p}$ is a step-$p$ adaptive weight. Following prior work [20, 43], to effectively encourage sparsity, $L_{Reg}$ always uses the $L_{1}$ regularization:
$\mathcal{L}_{Reg}=\left\|\frac{1}{n}(\boldsymbol{W_{r}}\boldsymbol{x_{t}})\right\|,$ (8)
$\beta_{p+1}=\beta_{p}\cdot e^{k\cdot(T_{s}-R_{s})},$ (9)
where $R_{s}$ is the current sparsity and $R_{s}=1-\frac{1}{n}(\boldsymbol{W_{r}}\boldsymbol{x_{t}})$, $k$ is a scaling factor to control the change of the step-wise weight. Such a design will make the training process more flexible, more stable, and with less extreme loss.
However, $L_{Reg}$ just focuses on increasing the sparsity, forcing the model to be more sparsely activated, which usually makes all output prone to 00. Considering this, we design a Head Activation Enhancement (HAE) Loss to activate more heads to avoid some experts always being asleep:
$\mathcal{L}_{HAE}=e^{2\cdot(R_{s}-T_{s})}-1\quad\text{if }R_{s}>T_{s}.$ (10)
$\mathcal{L}_{HAE}$ helps some heads to be more active when they fall into a state of zero. Combining the SPR and HAE, a balanced activation ratio can be guaranteed. Our final training objective is designed with the language-model used cross-entropy loss $\mathcal{L}_{LM}$:
$\mathcal{L}=\mathcal{L}_{LM}+\mathcal{L}_{SPR}+\mathcal{L}_{HAE}.$ (11)
3.4 Training Pipeline
We implement a two-stage training strategy, transitioning from image-summary pretraining to video-QA fine-tuning, teaching the model to progressively learn both static and dynamic subject-specific features. This strategy adapts our PVChat model from static image understanding to dynamic video reasoning. The architecture is shown in Fig. 4(a).
Stage 1: Image Understanding. In the first stage, the model focuses on learning static subject-specific knowledge from separate images, where we extract the first frame of each video as input. To preserve the pre-trained visual encoder’s capability, we freeze the visual encoder and only train the ReMoH components, by employing Low-Rank Adaptation (LoRA) [10] to efficiently fine-tune the Mistral-7B-Instruct-v0.3 language model [11], which significantly reduces trainable parameters while maintaining performance.
The QA pairs used in this stage primarily consist of existence verification (e.g., “Is <sks>visible in this video?”) and attribute description questions (e.g.,“What is <sks>wearing in this video?”). These questions guide the model to focus on static subject-specific features, which helps the model to do personalized image chat.
Stage 2: Video Reasoning. In the second stage, we extend the model to capture temporal dynamics and environmental interactions. We selectively train several later blocks in the visual encoder to enhance cross-frame feature integration in the personalized scenarios. This modification enables the model to recognize how subjects interact with their environments and what the subject is doing over time.
Here, we expand our QA dataset with action recognition questions (e.g., “What movements or actions does <sks>perform here?”) and location identification questions (e.g., “Can you describe <sks>’s location relative to others?”). Utilizing both positive and negative samples for training, our PVChat model becomes robust and precise.
4 Experiments
Figure 5: Examples of PVChat’s ability with a learned video (e.g., a man named <Sh>and another man named <Ho>). PVChat can recognize and answer questions about the personalized concept in various scenarios, such as medical scenarios (left) and TV series (right).
4.1 Settings
[…]
Datasets. To establish the dataset, we collect six different scenarios, including (Friends(6), Good Doctor(5), Ne Zha(2), doctor(3), patient(3), Big Bang(6)) $25$ characters (including doctor, patient, cartoon character, engineer, professor, etc.), and collect $304$ original videos and $2{,}304$ extended videos with over $30{,}000$ QA pairs. $300$ prompt descriptions for different scenarios are designed for different ages and genders. The structure and distribution of our dataset can be found in Fig. 6. For each character, we randomly select one video for the training set as our refer video and the other for the evaluation set as query video. We show more details and examples in our supplementary.
Evaluation Metrics across Multiple Datasets. Following the evaluation of LLaVA [25] and Yo’LLaVA [34], we rigorously assess PVChat with state-of-the-art ViLLMs using five complementary metrics across four question categories: existence, appearance, action, and location queries. The final results presented represent the averaged performance across all datasets, offering a holistic view of model capabilities.
Our evaluation metrics incorporate: (1) Accuracy (Acc) to measure the correctness of the presence or absence of subjects; (2) BLEU and BERTScore (BS) to quantify the textual similarity between model’s response and the ground truth; (3) Entity Specificity (ES) to assess if responses contain personalized information; (4) Descriptive Completeness (DC) to measure the quality of details inside responses. More details about the evaluation metrics can be found in the supplementary. For appearance, action, and location queries, we employ BLEU, BS, ES, and DC to comprehensively evaluate both semantic accuracy and conversational quality of generated responses.
4.2 Comparison with State-of-the-Art Models
Quantitative Results. To verify PVChat’s effectiveness, we present the quantitative results in Tab. 1, which is compared with state-of-the-art ViLLMs, including InternVideo2 [50], VideoLLaMA2 [5].
The result demonstrates that our PVChat exhibits a strong ability for personalized video understanding, offering high-accuracy and comprehensive responses. Especially for the negative sample, another model always replies that the object character is existent and no real understanding of the characteristic of the object. Since these models do not support multi-video input in a single-round conversation, we are not able to input the reference video and query video simultaneously. To ensure fairness, we find that InternVideo2 [50] supports the multi-round conversation, allowing us to use the reference video in the first round conversation, and input our query video in the second round. VideoLLaMA2 [5] does not support the multi-round conversation, so we input a detailed description of the reference video, and then test the query video using the character’s detailed description.
[…]
InternVideo2 [50] 0.342 0.046 0.875 3.041 1.812
VideoLLaMA2 [5] 0.470 0.082 0.890 3.012 3.301
PVChat (Ours) 0.901 0.562 0.952 4.940 4.201
Table 1: Quantitative evaluation of our method against state-of-the-art methods [50, 5]. Compared with these SOTA models, our model exhibits superior performance across five metrics.
Figure 6: The hierarchical structure of our prompt library, which is carefully divided into four levels, such as gender, age, and scenarios, and provides different descriptions according to the specific subject.
Qualitative Analysis. We conduct qualitative assessments on some challenging scenarios, including single-subject and multi-subject settings in various interactive environments and actions in Fig. 5. The result shows that our PVChat outperforms other models [50, 5] in different settings, demonstrating our superior personalized video understanding ability. Due to the shortage of personalized training and subject-specific data, previous models cannot concisely capture personalized information and show reasonable responses with correct identity.
Figure 7: The comparison of expert heads activation between MoH [12] and ReMoH in different layers, where $\text{H}_{i}$ represents the $i^{th}$ head. Orange refers to the video without the target individual, while blue represents the video having the character.
4.3 Ablation Study
Analysis of ReMoH. To explore the impact of ReMoH, we conduct experiments as shown in Tab. 2 and visualize the heads of expert activation ratio in Fig. 7. We choose the typical transformer architecture Q-former [57] with the LLM backbone [11] as our baseline. Compared with the baseline and incorporating the typical MoH [12], our ReMoH results in respective improvements of $18.6\%$, $9.8\%$ on the Acc, where our ReMoH significantly enhances the model’s capability of personalizing understanding and chatting.
In Fig. 7, we find that ReMoH effectively allocates some certain heads of expert to extract the video feature about the target character, while MoH [12] almost evenly distributes expert heads to learn the knowledge. For example, if using MoH [12], in most layers, no matter the character appears in the video or not, the first and second heads are always activated for video reasoning, verifying that these heads of expert actually do not learn about specific information. On the contrary, with the utility of ReMoH, in most layers it is observed that the activation of expert heads undergoes significant changes from no characters to having characters. This pattern supports that the model can learn personalized knowledge with ReMoH better, improving the QA performance related to specific identity, as presented in Tab. 2.
[…]
Baseline 0.733 0.550 0.904 4.735 4.142
Baseline + MoH [12] 0.813 0.558 0.926 4.939 4.191
Baseline + ReMoH 0.901 0.562 0.952 4.940 4.201
Table 2: Ablations on the proposed ReMoH, where ReMoH significantly outperforms MoH in Acc.
Influence of SPR and HAE. We conducted ablations to verify the effectiveness of SPR and HAE losses. As shown in Tab. 3, the introduction of SPR successfully smooths loss fluctuations and enhances training stability compared with MoH [12]. However, models with only SPR suffer from very low expert head activation ratio, limiting personalized feature extraction. Adding HAE loss alongside SPR, the head activation diversity increases significantly, allowing the model to capture more nuanced personalized representations.
[…]
PVChat w/o SPR and MAE nan
PVChat w/o MAE 0.217 0.085 0.746 0.555 0.926 4.913 4.112
PVChat 0.552 0.028 0.901 0.562 0.952 4.940 4.201
Table 3: Ablations on SPR and HAE losses. It verifies that SPR and HAE guarantee stability and enhance learning of the expert heads.
Contribution of Each Type of Data. We conduct an ablation on the dataset creation. Tab. 4 presents all metrics for the personalization conversion. Using vanilla datasets (with only one positive sample) fails to perform well on personalized questions, always responding with “yes” to all questions. After injecting the negative samples, it could recognize the identity of the character to a certain degree, but still suffers from some complex questions about action and location. After training with the generated data from ConsisID [56] and LivePortrait [9], the model is capable of understanding the specific information more accurately and robustly.
[…]
one positive 0.464 0.417 0.905 4.826 3.947
+Negative 0.584 0.418 0.931 4.899 4.120
+ConsisID Positive 0.781 0.532 0.927 4.929 4.132
+LivePortrait Positive 0.901 0.562 0.952 4.940 4.201
Table 4: Ablations on the dataset collection, where combining all types of designed data, the model performs accurately and robustly.
Influence of Token Numbers. We set each question prompt corresponding to $N=16$ tokens per character. As the token number increases, the metrics overall increase, but decrease when larger than 16. This is because too many tokens make the model hard to capture the characteristics of the subject. The detailed results can be found in our supplementary.
5 Conclusion
This paper introduces PVChat, the first personalized ViLLM that enables personalized subject learning from a single video. We collect comprehensive datasets over $2{,}300$ videos, propose ReMoH for specific knowledge learning, and employ a two-stage training strategy from image pre-training to video fine-tuning, with SPR and HAE losses. Evaluations show the superiority of our model in personalized video QA, making it potential for smart healthcare and home scenarios.
Supplementary Material:
PVChat: Personalized Video Chat with One-Shot Learning
In this supplemental material, the readers can find:
  • Experimental details about the ablation study of token number;
  • Details about multiple metrics utilized for personalized QA quality;
  • More examples of recognition question answering;
  • More experiment detail about Multi Character training
  • Some examples of existential questions and answers for single and two entities;
  • Presentation of templates for GPT prompt queries;
  • Presentation of 25 characters;
S1 Ablation Study About the Token Number
We set each question prompt corresponding to $N=16$ tokens per character. If $N=0$, training is only with the <sks>token. As the token number increases, the metrics overall increase. However, when the number is larger than $16$, the performance decreases. This is because too many tokens can make the model hard to capture the characteristics of the subject. The detailed experimental results are shown in Tab. 1.
[…]
0 0.801 0.495 0.939 5.00 4.00
4 0.871 0.592 0.951 5.00 4.50
8 0.890 0.553 0.949 5.00 4.58
12 0.895 0.564 0.949 5.00 4.41
16 0.922 0.606 0.952 4.74 4.38
20 0.882 0.554 0.949 4.96 4.25
Table 1: Ablations on the number of tokens in <Aa>.
S2 Details of the Evaluation Metrics
Here we add more details of the metrics which are utilized to measure the quality of personalized video chat:
(1) Accuracy: Specifically designed to evaluate binary existence questions, measuring the model’s ability to correctly identify the presence or absence of objects.
(2) BLEU and BERTScore: These metrics quantify the textual similarity between generated responses and ground truth answers, capturing linguistic precision and semantic alignment.
(3) Entity Specificity (ES): Evaluated on a 1-5 scale, this metric assesses whether responses contain personalized, contextually relevant details rather than generic statements.
(4) Descriptive Completeness (DC): Also rated on a 1-5 scale, DC measures the logical coherence, factual correctness, and comprehensive nature of responses. This metric evaluates whether answers are thoroughly developed with appropriate supporting details and proper reasoning.
S3 More Examples of Our PVChat Model
From Fig. 1 to Fig. 5, more personalized video chat examples of various scenarios are displayed. For example, in Fig. 1, there are three characters named <Cl>, <Ja>and <Xo>in a laboratory scene, which is quite challenging for video understanding. Our PVChat not only accurately identifies these three men by their names but also gives reasonable suggestions for some additional questions about giving gifts. What’s more, when the model is asked about their behavior, it successfully captures their locations and can do reasoning to guess this is a professional setting. This demonstrates that our model has strong personalized reasoning capability, even in a multiple subject setting. More details can be seen in the figures.
Figure 1: Example of PVChat.
Figure 2: Example of PVChat.
Figure 3: Example of PVChat.
Figure 4: Example of PVChat.
Figure 5: Example of PVChat.
S4 More experiment detail about Multi-Character training
Epoch One person training needs 8 epochs, and two people training need 16 epochs, and three people training need 24 epochs.
[…]
Crucially, these query tokens are associated with corresponding masks that are only deactivated during the training of their respective subject videos. This selective masking strategy ensures that each query embedding specifically attends to its designated subject, effectively preventing cross-contamination between different identity representations.
[…]
S5 Example of 4 Different Questions
From Tab. 2 to Tab. 5, the detailed existence question for one entity and two entities are displayed. We show the other three questions about the appearance, action, and location in Tab. 6.
Index Question Yes Answer
1 Is there any trace of <sks> in this footage? Yes, <sks> is in this video.
2 Can you detect <sks> in this video clip? I can confirm that <sks> appears.
3 Does <sks> show up anywhere in this recording? <sks> is present in this recording.
4 Is <sks> visible in this video? The video contains <sks>.
5 Could you verify if <sks> is here? I’ve identified <sks>.
6 Does this footage include <sks>? <sks> is shown in this video.
7 Can you spot <sks> in this clip? Yes, <sks> appears here.
8 Is <sks> present in this video? I can verify that <sks> is present.
9 Does <sks> appear in this footage? The footage shows <sks>.
10 Can you tell if <sks> is shown here? <sks> is in this video clip.
11 Is <sks> in this video segment? I’ve detected <sks>.
12 Can you confirm <sks>’s presence? Yes, <sks> is featured.
13 Does this clip contain <sks>? The video includes <sks>.
14 Is <sks> featured in this recording? I can see <sks>.
15 Can you find <sks> in this video? <sks> is definitely here.
16 Is <sks> shown in any frame? Yes, I’ve found <sks>.
17 Does this video show <sks>? This video shows <sks>.
18 Is <sks> visible anywhere? <sks> is visible.
19 Can you see <sks>? Yes, <sks> has been captured.
20 Is <sks> in this video? The video clearly shows <sks>.
21 Can you recognize <sks>? I’ve spotted <sks>.
22 Does <sks> appear at all? <sks> appears in this video.
23 Is <sks> recorded here? Yes, this footage contains <sks>.
24 Can you identify <sks>? I can recognize <sks>.
25 Is <sks> present? <sks> is clearly visible.
Table 2: Example of positive existence verification answering (Single entity).
Index Question No Answer
1 Is there any trace of <sks> in this footage? No, <sks> is not in this video.
2 Can you detect <sks> in this video clip? I cannot detect <sks>.
3 Does <sks> show up anywhere in this recording? This video does not contain <sks>.
4 Is <sks> visible in this video? <sks> is not shown.
5 Could you verify if <sks> is here? There is no sign of <sks>.
6 Does this footage include <sks>? <sks> does not appear.
7 Can you spot <sks> in this clip? I can confirm <sks> is not here.
8 Is <sks> present in this video? The footage does not include <sks>.
9 Does <sks> appear in this footage? There’s no evidence of <sks>.
10 Can you tell if <sks> is shown here? <sks> is not in this video.
11 Is <sks> in this video segment? I’ve checked, <sks> is not present.
12 Can you confirm <sks>’s presence? This video does not show <sks>.
13 Does this clip contain <sks>? I see no sign of <sks>.
14 Is <sks> featured in this recording? <sks> is absent.
15 Can you find <sks> in this video? The video does not show <sks>.
16 Is <sks> shown in any frame? I cannot find <sks>.
17 Does this video show <sks>? <sks> is not visible.
18 Is <sks> visible anywhere? I can verify <sks> is not here.
19 Can you see <sks>? The video has no <sks>.
20 Is <sks> in this video? <sks> does not exist in this video.
21 Can you recognize <sks>? I find no trace of <sks>.
22 Does <sks> appear at all? This clip does not contain <sks>.
23 Is <sks> recorded here? <sks> is not present.
24 Can you identify <sks>? I cannot identify <sks>.
25 Is <sks> present? There is no <sks> here.
Table 3: Example of negative existence verification answering (Single entity).
Index Question Both Yes Answer
1 Is there any trace of <sks1> or <sks2> in this footage? Both <sks1> and <sks2> are present in this video.
2 Can you detect <sks1> or <sks2> in this video clip? I can detect both <sks1> and <sks2> in the footage.
3 Do <sks1> or <sks2> show up anywhere in this recording? The video shows both <sks1> and <sks2> clearly.
4 Are <sks1> or <sks2> visible in this video? <sks1> and <sks2> are both visible in this recording.
5 Could you verify if <sks1> or <sks2> are here? I’ve identified both <sks1> and <sks2> in the clip.
6 Does this footage include <sks1> or <sks2>? Both <sks1> and <sks2> appear in this video.
7 Can you spot <sks1> or <sks2> in this clip? The footage contains both <sks1> and <sks2>.
8 Are <sks1> or <sks2> present in this video? I can see both <sks1> and <sks2> in frame.
9 Do <sks1> or <sks2> appear in this footage? <sks1> and <sks2> are both featured in this video.
10 Can you tell if <sks1> or <sks2> are shown here? The recording shows both <sks1> and <sks2> present.
11 Are <sks1> or <sks2> in this video segment? I’ve spotted both <sks1> and <sks2> in the footage.
12 Can you confirm <sks1> or <sks2>’s presence? Both <sks1> and <sks2> are captured in this clip.
13 Does this clip contain <sks1> or <sks2>? The video includes both <sks1> and <sks2>.
14 Are <sks1> or <sks2> featured in this recording? I can confirm the presence of both <sks1> and <sks2>.
15 Can you find <sks1> or <sks2> in this video? <sks1> and <sks2> are both shown in the recording.
16 Are <sks1> or <sks2> shown in any frame? Both figures, <sks1> and <sks2>, are visible.
17 Does this video show <sks1> or <sks2>? I’ve found both <sks1> and <sks2> in the video.
18 Are <sks1> or <sks2> visible anywhere? The footage displays both <sks1> and <sks2>.
19 Can you see <sks1> or <sks2>? Both <sks1> and <sks2> are identifiable here.
20 Are <sks1> or <sks2> in this video? I can recognize both <sks1> and <sks2>.
21 Can you recognize <sks1> or <sks2>? <sks1> and <sks2> both appear in this recording.
22 Do <sks1> or <sks2> appear at all? The video features both <sks1> and <sks2>.
23 Are <sks1> or <sks2> recorded here? Both <sks1> and <sks2> are clearly visible.
24 Can you identify <sks1> or <sks2>? I’ve detected the presence of both <sks1> and <sks2>.
25 Are <sks1> or <sks2> present? The clip shows both <sks1> and <sks2>.
Table 4: Example of positive existence verification question answering (Two entities).
Index sks1 Yes, sks2 No Both No
1 Is there any trace of <sks1> or <sks2> in this footage? I can confirm that <sks1> appears, but <sks2> is not present.
Neither <sks1> nor <sks2> appear in this video.
2 Can you detect <sks1> or <sks2> in this video clip? The video shows <sks1>, though there’s no sign of <sks2>.
I cannot detect either <sks1> or <sks2>.
3 Do <sks1> or <sks2> show up anywhere in this recording? <sks1> is visible, but <sks2> is absent.
The video contains neither <sks1> nor <sks2>.
4 Are <sks1> or <sks2> visible in this video? I’ve detected <sks1>, while <sks2> does not appear.
Both <sks1> and <sks2> are absent.
5 Could you verify if <sks1> or <sks2> are here? The video contains <sks1>, but <sks2> is not shown.
There is no sign of either <sks1> or <sks2>.
6 Does this footage include <sks1> or <sks2>? <sks1> is present, however <sks2> is not in this clip.
Neither <sks1> nor <sks2> are shown.
7 Can you spot <sks1> or <sks2> in this clip? I can see <sks1>, but there’s no trace of <sks2>.
I confirm both <sks1> and <sks2> are not present.
8 Are <sks1> or <sks2> present in this video? The footage includes <sks1>, though <sks2> is not visible.
The footage does not include <sks1> or <sks2>.
9 Do <sks1> or <sks2> appear in this footage? <sks1> appears, but <sks2> is not featured.
There’s no evidence of either <sks1> or <sks2>.
10 Can you tell if <sks1> or <sks2> are shown here? I’ve spotted <sks1>, while <sks2> is nowhere to be seen.
Neither <sks1> nor <sks2> are visible.
11 Are <sks1> or <sks2> in this video segment? <sks1> is clearly visible, but <sks2> is not.
I’ve checked, both <sks1> and <sks2> are absent.
12 Can you confirm <sks1> or <sks2>’s presence? The recording shows <sks1>, though <sks2> is absent.
This video shows neither <sks1> nor <sks2>.
13 Does this clip contain <sks1> or <sks2>? I can identify <sks1>, but <sks2> doesn’t appear.
I see no sign of <sks1> or <sks2>.
14 Are <sks1> or <sks2> featured in this recording? <sks1> is present, while <sks2> is not.
Both <sks1> and <sks2> are not in the recording.
15 Can you find <sks1> or <sks2> in this video? The clip features <sks1>, but there’s no sign of <sks2>.
The video does not contain <sks1> or <sks2>.
Table 5: Example of mixed existence verification question answering (Two entities).
Index Question
1 What activity is <sks> engaged in during this video?
2 Could you describe what <sks> is doing in this footage?
3 What specific actions can you observe <sks> performing in this recording?
4 What movements or actions does <sks> perform here?
5 Can you describe <sks>’s behavior in this sequence?
6 What is <sks> wearing in this video?
7 Could you describe <sks>’s outfit in this footage?
8 What color and style of clothing is <sks> dressed in?
9 How would you describe <sks>’s appearance and attire?
10 What notable features can you see in <sks>’s clothing?
11 Where is <sks> positioned in this video?
12 What color and style of clothing is <sks> dressed in?
13 Can you describe <sks>’s location relative to others?
14 Which part of the scene does <sks> appear in?
15 How does <sks>’s position change throughout the video?
16 Where can <sks> be found in this footage?
Table 6: Example of negative existence verification question answering (Single entity).
S6 Example of queries prompt for GPT or Internvideo
As shown in Fig. 7, this prompt will be used to get the age and gender of the character in the video by the video understanding model, and the age and gender information will be used in the following steps. And for the general answer from the video, understanding model will use the second one prompt to query ChatGPT to capture the general description like the human, he, she, and this will be replaced by the specific <sks>and then to construct our personalized QA pairs.
As shown in Fig. 7, leveraging the character’s age and gender attributes extracted in the previous stage, we enrich these core identity markers with detailed descriptors to create comprehensive prompts for (image+text)-to-image models. This approach significantly enhances identity consistency across generated images, maintaining coherent character representation while accommodating diverse poses and contexts. The resulting outputs demonstrate superior identity preservation compared to methods lacking such demographic anchoring.
S7 Presentation of 25 characters
As shown in Fig. 8, our dataset is composed of Friends(6), Good Doctor(5), Ne Zha(2), doctor(3), patient(3), Big Bang(6)) 25 characters.
Figure 6: Prompt for Internvideo and GPT query
Figure 7: Prompt for Photomaker synthetic the ID consist photo
Figure 8: Display of all dataset.

### v4

@@
 We evaluate PVChat on diverse datasets covering medical scenarios, TV series, anime, and real-world footage, demonstrating its superiority in personalized feature understanding after learning from a single video, compared to state-of-the-art ViLLMs.
+Code is at PVChat.
 Figure 1: Examples of PVChat’s ability with one-shot learning (e.g., <Nz>and <Ab>). PVChat can answer questions about the personalized information correctly while other models [5, 50] fail.
@@
 Recent advancements in Video Large Language Models (ViLLMs) have showcased impressive capabilities across a wide range of tasks, e.g., video captioning, temporal action localization, and video question answering [5, 50, 31, 49].
-These models, trained on vast datasets, exhibit broad knowledge across multiple domains, such as medical assistance [17] and autonomous driving [55], significantly extending the functional scope of ViLLMs.
+These models, trained on vast datasets, exhibit broad knowledge across multiple domains, such as medical assistance [17] and autonomous driving [55, 56], significantly extending the functional scope of ViLLMs.
 Despite these achievements, current ViLLMs are largely confined to general-purpose content understanding and struggle in personalized scenarios that require distinguishing specific individuals.
@@
 The key is to exploit off-the-shelf toolboxes to generate identity-preserving positives and synthesize hard negatives for confusingly similar faces.
-To support the research, we construct a diverse dataset containing $6$ individual scenarios, $304$ original videos, $2{,}304$ extended generated videos with over $30{,}000$ QA pairs, which will be publicly released to facilitate future research.
+To support the research, we construct a diverse dataset containing $6$ individual scenarios, $304$ original videos, 2,3042{,}304 extended generated videos with over 30,00030{,}000 QA pairs, which will be publicly released to facilitate future research.
   • We design the ReLU Routing Mixture-of-Heads module to enhance the effective extraction of individual-specific characteristics from videos.
@@
 3.
-Through this comprehensive pipeline, each input natural video is transformed into $81$ different videos accompanied by $1{,}455$ question-answer pairs, significantly expanding our training dataset while maintaining subject identity consistency and guaranteeing the model’s robustness.
+Through this comprehensive pipeline, each input natural video is transformed into $81$ different videos accompanied by 1,4551{,}455 question-answer pairs, significantly expanding our training dataset while maintaining subject identity consistency and guaranteeing the model’s robustness.
 3.3 ReLU Routing Mixture-of-Heads Attention
@@
 For $T_{1}$ input tokens $\boldsymbol{X}_{1}\in\mathbb{R}^{T_{1}\times d_{in}}$ where $d_{in}$ is the input dimension, and $T_{2}$ tokens $\boldsymbol{X}_{2}\in\mathbb{R}^{T_{2}\times d_{in}}$, attention is conducted as follows:
-[…]
+ $\begin{array}[]{l}\text{ Attention }(\boldsymbol{Q},\boldsymbol{K},\boldsymbol{V})=\operatorname{Softmax}\left(\frac{\boldsymbol{Q}\boldsymbol{K}^{\top}}{\sqrt{d_{k}}}\right)\boldsymbol{V},\\ \text{where}\>\boldsymbol{Q}=\boldsymbol{X}_{1}\boldsymbol{W}_{Q},\boldsymbol{K}=\boldsymbol{X}_{2}\boldsymbol{W}_{K},\boldsymbol{V}=\boldsymbol{X}_{2}\boldsymbol{W}_{V}.\end{array}$ (1)
 To enhance the representation of the attention layer, Vaswani [48] proposed to use multiple attention heads to operate on the projection of the input.
@@
 Specifically, the score for each head is defined as:
-$s_{i}=\left\{\begin{array}[]{ll}\alpha_{1},&\text{ if }1\leq i\leq n,\\ \alpha_{2}\operatorname{ReLU}\left(\boldsymbol{W}_{r}\boldsymbol{x}_{t}\right)_{i},&\text{ if }n<i\leq n+m,\end{array}\right.$ (5)
+$s_{i}=\left\{\begin{array}[]{ll}\alpha_{1},&\text{ if }1\leq i\leq n,\\ \alpha_{2}\operatorname{ReLU}\left(\boldsymbol{W}_{r}\boldsymbol{x}_{t}\right)_{i},&\text{ if }n<i\leq n+m,\end{array}\right$.
+(5)
 where $n$ and $m$ are the number of shared heads and routed heads, respectively.
@@
 Such a design will make the training process more flexible, more stable, and with less extreme loss.
-However, $L_{Reg}$ just focuses on increasing the sparsity, forcing the model to be more sparsely activated, which usually makes all output prone to 00.
+However, $L_{Reg}$ just focuses on increasing the sparsity, forcing the model to be more sparsely activated, which usually makes all output prone to 0.
 Considering this, we design a Head Activation Enhancement (HAE) Loss to activate more heads to avoid some experts always being asleep:
@@
 4.1 Settings
-[…]
+Implementation Details.
+Our LLM backbone is based on the Mistral-7B-Instruct-v0.3 [11].
+All models are trained on one NVIDIA L20 GPU.
+The first stage of training takes one epoch, and the second stage takes seven epochs.
+We set the batch size to $2$, the learning rates of token embedding, ReMoH, LLM Lora to $\times 10^{-4}, 1×1$$\times 10^{-5}$, and $1\times 10^{-5}$, training time needs $3h$, video resolution is $1080\times 1920$.For visual encoder each video will evenly sample 8 frames.
 Datasets.
-To establish the dataset, we collect six different scenarios, including (Friends(6), Good Doctor(5), Ne Zha(2), doctor(3), patient(3), Big Bang(6)) $25$ characters (including doctor, patient, cartoon character, engineer, professor, etc.), and collect $304$ original videos and $2{,}304$ extended videos with over $30{,}000$ QA pairs.
+To establish the dataset, we collect six different scenarios, including (Friends(6), Good Doctor(5), Ne Zha(2), doctor(3), patient(3), Big Bang(6)) $25$ characters (including doctor, patient, cartoon character, engineer, professor, etc.), and collect $304$ original videos and 2,3042{,}304 extended videos with over 30,00030{,}000 QA pairs.
 $300$ prompt descriptions for different scenarios are designed for different ages and genders.
@@
 This paper introduces PVChat, the first personalized ViLLM that enables personalized subject learning from a single video.
-We collect comprehensive datasets over $2{,}300$ videos, propose ReMoH for specific knowledge learning, and employ a two-stage training strategy from image pre-training to video fine-tuning, with SPR and HAE losses.
+We collect comprehensive datasets over 2,3002{,}300 videos, propose ReMoH for specific knowledge learning, and employ a two-stage training strategy from image pre-training to video fine-tuning, with SPR and HAE losses.
 Evaluations show the superiority of our model in personalized video QA, making it potential for smart healthcare and home scenarios.
@@
 Epoch One person training needs 8 epochs, and two people training need 16 epochs, and three people training need 24 epochs.
-[…]
+ReMoH Within the ReMoH architecture, we implement cross-attention mechanisms at alternating layers to facilitate interaction between video embeddings (extracted by the visual encoder) and query embeddings.
+Our approach adaptively scales with the number of subjects: for single-person training, we incorporate an additional $16\times\text{token\_length}\times 768$ dimensional feature space; for two-person training, this expands to $32\times\text{token\_length}\times 768$ and for three-person training, it further increases to $48\times\text{token\_length}\times 768$.
 Crucially, these query tokens are associated with corresponding masks that are only deactivated during the training of their respective subject videos.